Solution and Answer please, Thank you!
- If 1000 articles of a give type can be turned out by a first machine in 8 hours, by a second in 5 hours, and by a third in 4 hours, how long will it take to turn out the articles with all machines working?

Answer by richwmiller(17219)

a=8 b=5 c=4
1/a+1/b+1/c=1/x
1/8+1/5+1/4=1/x
5/40+8/40+10/40=1/x
23/40=x
40/23=x
x=1.73913043 hours working together
check
1/a+1/b+1/c=1/x
1/8+1/5+1/4=1/40/23
5/40+8/40+10/40=23/40
23/40=23/40
ok